# Large-scale moral machine experiment on large language models

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) and their potential integration into autonomous driving systems necessitates understanding their moral decision-making capabilities. While our previous study examined four prominent LLMs using the Moral Machine experimental framework, the dynamic landscape of LLM development demands a more comprehensive analysis. Here, we evaluate moral judgments across 51 different LLMs, including multiple versions of proprietary models (GPT, Claude, Gemini) and open-source alternatives (Llama, Gemma), to assess their alignment with human moral preferences in autonomous driving scenarios. Using a conjoint analysis framework, we evaluated how closely LLM responses aligned with human preferences in ethical dilemmas and examined the effects of model size, updates, and architecture. Results showed that proprietary models and open-source models exceeding 10 billion parameters demonstrated relatively close alignment with human judgments, with a significant negative correlation between model size and distance from human judgments in open-source models. However, model updates did not consistently improve alignment with human preferences, and many LLMs showed excessive emphasis on specific ethical principles. These findings suggest that while increasing model size may naturally lead to more human-like moral judgments, practical implementation in autonomous driving systems requires careful consideration of the trade-off between judgment quality and computational efficiency. Our comprehensive analysis provides crucial insights for the ethical design of autonomous systems and highlights the importance of considering cultural contexts in AI moral decision-making.
@@ Introduction
The rapid advancement of Large Language Models (LLMs) has inaugurated a new era in artificial intelligence (AI), demonstrating unprecedented capabilities across diverse domains [1, 2, 3]. These models, exemplified by ChatGPT [4, 5] (developed by OpenAI) and its counterparts, exhibit remarkable proficiency in natural language processing, content generation, complex problem-solving, and decision-making. As LLMs evolve, their potential integration into critical systems has become a subject of intense scrutiny and research [6, 7, 8].
Autonomous driving represents a field where LLM application holds particular promise while raising significant ethical questions [9, 10, 11, 12]. The automotive industry is actively exploring LLM integration to enhance autonomous vehicles’ capabilities, potentially improving environmental comprehension, complex scenario interpretation, nuanced decision-making, and communication with passengers and other road users [13, 14, 15, 16, 17, 18, 19]. This integration spans various aspects of autonomous driving, including perception, navigation, planning, and control.
The Moral Machine experiment [10] emphasized the importance of ethical considerations in AI decision-making for autonomous systems. This study presented participants with scenarios involving unavoidable accidents to assess public opinion on how autonomous vehicles should act in morally challenging situations. The experiment provided a framework for exploring human moral preferences, identifying trends such as prioritizing human lives over animals and favoring the preservation of a greater number of lives. While acknowledging its limitations [20, 21, 22, 23, 24, 25], this work has stimulated extensive research into the ethical implications of AI in autonomous systems [26, 27, 28, 29].
@@
Following the methodology established in our previous study [33], we generated a comprehensive set of Moral Machine scenarios to evaluate LLM ethical decision-making. These scenarios presented dilemmas involving an autonomous vehicle facing an unavoidable accident, requiring a choice between two outcomes (“Case 1” and “Case 2”) with varying ethical implications. The scenarios were designed through a process of constrained randomization, exploring six primary dimensions: species (humans vs. pets), social value (higher vs. lower perceived social status), gender (female vs. male), age (younger vs. older), fitness (physically favored vs. less fit individuals), and utilitarianism (smaller vs. larger group). Additionally, three secondary dimensions were incorporated: interventionism (swerving vs. continuing straight), relationship to the autonomous vehicle (passengers vs. pedestrians), and legal considerations (e.g., adherence to traffic signals). The scenarios were carefully constructed to balance complexity and clarity, ensuring that they effectively probed the ethical reasoning capabilities of the LLMs under study.
@@
This study expanded upon our previous study [33] by incorporating a comprehensive set of large language models (LLMs), including the latest versions available. These LLMs were utilized to generate responses to moral machine scenarios.
@@
In the open-source domain, we included Meta’s Llama family [36], comprising Llama 3 and Llama 3.1 in 8B and 70B versions, as well as Llama 3.2 in 1B and 3B versions. A comprehensive evaluation of Google’s Gemma model family [37, 42] was conducted, encompassing multiple versions and scales from 2B-it to 27B-it, including the derivative DataGemma RIG 27B-it model [43]. Additional open-source models evaluated in this study included Large Model Systems Organization’s Vicuna v1.5 (7B and 13B) [44], Mistral (7B Instruct v0.2) [45], Mistral-NeMo [46], CohereForAI’s Command R+ [47], and Microsoft’s Phi-3.5 [48] (both MoE-instruct and mini-instruct versions).
@@
All models were used with their default parameter settings. The input prompts were consistent with those used in [33], requiring each model to select either “Case 1” or “Case 2” for each scenario.
@@
For analyzing LLM responses to the Moral Machine scenarios, we employed the conjoint analysis framework as described in [49] and implemented in the original study [10] on the Moral Machine experiment and our previous study [33]. This approach allows for a robust, non-parametric identification of causal effects without specific modeling assumptions.
@@ Results
Figure 1 presents radar charts characterizing the moral judgment tendencies of various LLMs across nine preferences quantified by AMCE values (see S1 Table. for exact numerical data). These charts compare the preferences across multiple versions and model sizes of GPT-3.5, GPT-4, GPT-4o/o1, Claude, Gemini, Llama, Gemma, and other LLMs against human preferences (see S1 Fig. for detailed differences between each model and human values).
Analysis revealed that many LLMs demonstrated large positive values (¿0.5) for No. Characters and Species dimensions, indicating tendencies similar to human judgments in prioritizing the preservation of more lives and favoring humans over animals. However, some models exhibited either inverse priorities compared to humans or disproportionately strong preferences in the same direction (differences in AMCE values approximately ¿0.3 compared to human values). We elaborate on these findings below, including the effects of model updates.
@@
Analysis of the GPT-3.5 family (Fig 1A) revealed interesting changes associated with model updates. While models from June 2023 to January 2024 exhibited similar preference patterns, the March 2023 version demonstrated distinctive judgment tendencies. For instance, regarding Law, the March version showed positive values favoring law-abiding individuals, aligning with human preferences but displaying stronger preferences than humans. However, subsequent versions shifted toward more neutral judgments with values approaching zero.
Similar changes were observed in Gender-related judgments. The March 2023 version showed strong positive values favoring females, but subsequent updates gradually reduced these values, converging toward human judgment levels. Regarding age-related decisions, while humans demonstrated a preference for younger individuals, all GPT-3.5 versions maintained neutral judgments with values near zero.
For Fitness-related judgments, the March and June 2023 versions showed negative values contrary to human preferences, prioritizing overweight individuals over physically fit ones. However, later versions gradually aligned more closely with human judgments. Regarding Relation to AV, all versions consistently showed positive values similar to humans, but with notably higher magnitudes, indicating a stronger preference for pedestrians over passengers. For Species-related judgments, while the March and June 2023 versions showed values comparable to humans, the November 2023 and January 2024 versions exhibited decreased values, maintaining but weakening the tendency to prioritize humans over pets.
@@
Analysis of the GPT-4 family (Fig 1B) revealed high consistency across models from March 2023 to January 2024, demonstrating that model updates had limited impact on moral judgments. Across all nine preference categories analyzed, the models exhibited directional alignment with human judgments, though showing stronger preferences in specific categories.
Most notably, in the Species and No. Characters categories, AMCE values approached 1.0, indicating nearly deterministic preferences for saving humans over animals and larger groups over smaller ones. Similarly, in the Gender, Law, and Relation to AV categories, while aligned directionally with human judgments, the models displayed larger AMCE values. Specifically, they showed more pronounced tendencies than humans in prioritizing females over males, law-abiding individuals over lawbreakers, and pedestrians over passengers.
Conversely, in the Fitness category, the models demonstrated trends diverging from human judgments. While humans showed a slight preference for physically fit individuals, the GPT-4 family exhibited a minor inverse preference, marginally favoring overweight individuals. However, this divergence was relatively minor compared to the pronounced preferences observed in other categories.
@@
Analysis of the GPT-4o family (Fig 1C) similarly demonstrated high consistency across models, with model updates showing limited impact on moral judgments. While this family exhibited AMCE values comparable to human judgments in the Gender and No. Characters categories, notable differences from human preferences were observed across many categories.
Particularly in the Relation to AV category, all models except GPT-4o-mini showed substantially higher values compared to even the GPT-4 family, demonstrating an extreme preference for pedestrians over passengers. The strong preference for law-abiding individuals persisted, and all GPT-4o models showed significantly stronger tendencies than humans in prioritizing humans over pets. Conversely, these models demonstrated relatively weak preferences for younger individuals compared to human age-related judgments, while maintaining a slight, albeit weak, preference for overweight individuals in Fitness-related decisions.
The o1 model generally exhibited judgment patterns similar to the GPT-4o family, though the o1-mini model showed distinctive variations. Notably, o1-mini demonstrated near-neutral preferences aligned with human judgments in the Relation to AV category. However, it recorded AMCE values approaching 1.0 in the No. Characters category, indicating a strong preference for saving larger groups of lives.
@@
Analysis of the Claude family (Fig 1D) revealed a distinct bifurcation in judgment tendencies across models. Specifically, two characteristic groups emerged: one comprising Claude 3 Sonnet, Opus, and 3.5 Haiku, and another consisting of Claude 3 Haiku and Claude 3.5 Sonnet.
Claude 3 Sonnet, Opus, and 3.5 Haiku demonstrated distinctive trends that diverged from human judgments across multiple categories. These models showed negative AMCE values in Relation to AV, Fitness, and Social Status categories, indicating inverse preferences compared to humans: prioritizing passengers over pedestrians, overweight over physically fit individuals, and lower over higher social status. While these models aligned directionally with humans in Gender (favoring females over males) and Species (prioritizing humans over pets) categories, the magnitude of their preferences substantially exceeded human values.
Conversely, Claude 3 Haiku and Claude 3.5 Sonnet exhibited different patterns, including neutral age-related judgments that diverged from human preferences. These models also demonstrated excessive preferences, particularly in Relation to AV and Law categories, showing substantially higher AMCE values than humans in prioritizing pedestrians over passengers and law-abiding over law-breaking individuals. Claude 3.5 Sonnet notably demonstrated pronounced tendencies in Species and No. Characters categories, strongly favoring humans over pets and larger over smaller groups.
An interesting finding regarding model evolution revealed that the Sonnet model’s update from 3 to 3.5 generally resulted in closer alignment with human judgments. However, no significant changes in judgment patterns were observed between the June and October versions of Claude 3.5 Sonnet. Conversely, the Haiku model’s update from 3 to 3.5 resulted in greater divergence from human judgments.
@@
Analysis of the Gemini family (Fig 1E) revealed high inter-model consistency across most categories, while demonstrating notable changes in judgment tendencies with updates, particularly in the Relation to AV category. Specifically, while the Gemini 1.0 Pro model showed negative AMCE values indicating preference for passengers over pedestrians, the Gemini 1.5 Pro family shifted to positive values aligned with human judgments, prioritizing pedestrians over passengers.
Similar update-driven changes were observed in the Gemini 1.5 Flash family. While the 1.5 Flash Preview and 1.5 Flash 001 versions prioritized passengers, the update to 1.5 Flash 002 reversed this preference to favor pedestrians. However, the degree of pedestrian preference in the latest versions of both Gemini 1.5 Pro and 1.5 Flash substantially exceeded human values.
Regarding gender-related judgments, only the latest Gemini 1.5 Pro 002 showed values comparable to human judgments, while other versions demonstrated pronounced preferences for females over males. In the Species category, all versions showed stronger preferences than humans for prioritizing humans over pets. Conversely, all models consistently demonstrated inverse tendencies to humans in Fitness-related decisions, showing negative AMCE values favoring overweight over physically fit individuals.
Figure 1E includes PaLM 2, Google’s previous proprietary model, for comparison. PaLM 2 exhibited a distinctive tendency contrary to human judgments, prioritizing smaller over larger groups. In contrast, the Gemini family, Google’s current flagship models, aligned with human judgments across all versions in prioritizing the preservation of more lives.
@@
Analysis of the Llama family (Fig 1F) demonstrated that all models except Llama 3.1 8B showed positive AMCE values similar to humans in the No. Characters category, indicating consistent preferences for saving larger groups of lives.
However, other categories exhibited varying judgment patterns across versions and model sizes. For gender-related decisions, while Llama 2 (7B), 3.2 1B, and 3.2 3B models showed preferences for males contrary to human judgments, other models aligned with humans in favoring females. In the Species category, Llama 3 70B, 3.1 8B, and 3.1 70B models demonstrated clear preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or slight preferences for pets. Similarly, regarding law compliance, Llama 3 8B, 3 70B, and 3.1 70B models demonstrated preferences for law-abiding individuals, while other models exhibited neutral preferences.
Notably, larger models (70B) showed relatively small variations in judgments between versions and demonstrated closer alignment with human preferences. This observation suggests that model size may influence the consistency of judgment patterns.
@@
Analysis of the Gemma family (Fig 1G) revealed characteristic consistency across several categories. All models demonstrated nearly neutral judgments regarding law compliance and showed consistent preferences for passengers over pedestrians. Notably, these consistent trends, particularly in the pedestrian-passenger prioritization, diverged from human judgments.
However, numerous other categories exhibited varying judgment patterns across models. In the No. Characters category, while Gemma 2 2B and 9B models showed preferences for smaller groups contrary to human judgments, other models aligned with humans in favoring the preservation of more lives. In the Species category, the Gemma 2 series (2B, 9B, 27B) demonstrated preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or preferences for pets. Regarding Social Status, only Gemma 7B and 1.1 7B aligned with human preferences favoring higher social status, while other models exhibited either neutral judgments or preferences for lower social status.
These variations in judgment consistency and alignment with human preferences suggest potential correlations with model versions and parameter sizes.
@@
Analysis of other LLMs (Fig 1H) generally showed trends similar to human judgments, though with distinctive differences in preference strengths and specific categories. Notably, all models except Vicuna demonstrated higher AMCE values than humans in the Gender category, showing particularly strong preferences for females over males. Additionally, while Mistral and Command R+ aligned with human preferences in favoring larger groups, their preference magnitudes were notably stronger than human values.
Inverse judgment tendencies compared to humans were characteristically observed in the Relation to AV and Fitness categories. Specifically, all models except Vicuna 13B showed negative AMCE values in these categories, demonstrating preferences contrary to human judgments by prioritizing passengers over pedestrians and overweight over physically fit individuals.
Vicuna (7B and 13B) and Phi 3.5 mini exhibited distinctive judgment patterns that generally diverged from human preferences. These models showed neutral judgments regarding law compliance and demonstrated slight preferences for smaller groups, contrary to human tendencies. Particularly, while Vicuna 7B showed neutral judgments in Gender, Age, and Species categories, Vicuna 13B demonstrated characteristic preferences for males over females, pets over humans, and older over younger individuals.
Significantly, different judgment tendencies were observed within the same model families, such as between Vicuna 7B and Vicuna 13B, or between Phi 3.5 mini and Phi 3.5 MoE. These results suggest potential influences of model size on ethical judgments.
@@
The GPT-4 family demonstrated the closest alignment with human judgments, with the January 2024 version showing the minimum distance of 0.6. The GPT-3.5 family also exhibited relatively close alignment, with distances ranging from 0.8 to 0.9. However, the newer GPT-4o/o1 family, except for o1-mini, showed a slight increase in distance from human judgments.
In the Claude family, while Claude 3 Haiku and Claude 3.5 Sonnet demonstrated relatively close alignment with human judgments (distances 0.9–1.0), Claude 3 Sonnet and Opus showed notably larger distances (1.2). The Gemini family similarly exhibited significant version-dependent variations, with Gemini 1.5 Preview (April 2024) showing the minimum distance of 0.9, while subsequent versions sometimes demonstrated larger distances of 1.1. Additionally, 1.5 Flash models showed greater distances compared to 1.5 Pro models.
Among open-source models, the Gemma family’s latest Gemma 2 2B and 9B models showed notably large distances from human judgments (2.0), with distances decreasing as model size increased. The Llama family demonstrated similar trends, with 70B models showing closer alignment to human judgments compared to smaller (1–8B) models.
Regarding other LLMs, Mistral-Nemo and Command R+ demonstrated relatively close alignment with human judgments (distances 0.8–0.9), while Vicuna and Phi 3.5 showed larger distances (1.0–1.6).

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) and their potential integration into autonomous driving systems necessitates understanding their moral decision-making capabilities. While our previous study examined four prominent LLMs using the Moral Machine experimental framework, the dynamic landscape of LLM development demands a more comprehensive analysis. Here, we evaluate moral judgments across 52 different LLMs, including multiple versions of proprietary models (GPT, Claude, Gemini) and open-source alternatives (Llama, Gemma), to assess their alignment with human moral preferences in autonomous driving scenarios. Using a conjoint analysis framework, we evaluated how closely LLM responses aligned with human preferences in ethical dilemmas and examined the effects of model size, updates, and architecture. Results showed that proprietary models and open-source models exceeding 10 billion parameters demonstrated relatively close alignment with human judgments, with a significant negative correlation between model size and distance from human judgments in open-source models. However, model updates did not consistently improve alignment with human preferences, and many LLMs showed excessive emphasis on specific ethical principles. These findings suggest that while increasing model size may naturally lead to more human-like moral judgments, practical implementation in autonomous driving systems requires careful consideration of the trade-off between judgment quality and computational efficiency. Our comprehensive analysis provides crucial insights for the ethical design of autonomous systems and highlights the importance of considering cultural contexts in AI moral decision-making.
@@ Introduction
The rapid advancement of Large Language Models (LLMs) has inaugurated a new era in artificial intelligence (AI), demonstrating unprecedented capabilities across diverse domains [1, 2, 3, 4]. These models, exemplified by ChatGPT [5, 6] (developed by OpenAI) and its counterparts, exhibit remarkable proficiency in natural language processing, content generation, complex problem-solving, and decision-making. As these models advance from laboratory experiments to real-world applications, understanding their decision-making processes, particularly in ethical contexts, becomes crucial for ensuring their responsible deployment. This has become a subject of intense scientific investigation [7, 8, 9].
Autonomous driving represents a field where LLM application necessitates particularly careful evaluation [10, 11, 12, 13]. When autonomous vehicles make decisions on public roads, their choices directly affect human safety and well-being. The automotive industry’s investigation of LLM integration for enhancing autonomous vehicles’ capabilities includes various aspects: analyzing complex road conditions, interpreting traffic situations, planning appropriate responses, and importantly, making ethical decisions in challenging scenarios [14, 15, 16, 17, 18, 19, 20]. This integration spans critical functions from environmental comprehension to decision making, highlighting the need for thorough understanding of these models’ capabilities and limitations.
The Moral Machine experiment [11] provides a systematic framework for evaluating ethical decision-making in autonomous systems. This study presented participants with scenarios involving unavoidable accidents to assess how autonomous vehicles should behave in morally challenging situations. The experiment revealed clear patterns in human moral preferences, such as prioritizing human lives over animals and favoring the preservation of a greater number of lives. While acknowledging its limitations including methodological constraints in scenario design, challenges in cultural representation, and gaps between theoretical choices and practical implementation [21, 22, 23, 24, 25, 26], this work has stimulated extensive research into the ethical implications of AI in autonomous systems [27, 28, 29, 30]. These foundational insights, despite their limitations which we explore in detail in our discussion, have provided crucial frameworks for examining ethical decision making in autonomous systems.
@@
To comprehensively evaluate how different LLMs approach moral decision-making in autonomous driving scenarios and assess the effects of model updates on these judgments, we generated a systematic set of test scenarios following our previous methodology. These scenarios presented dilemmas involving an autonomous vehicle facing an unavoidable accident, requiring a choice between two outcomes (“Case 1” and “Case 2”) with varying ethical implications.
The scenarios were designed through a process of constrained randomization, exploring six primary dimensions: species (humans vs. pets), social value (higher vs. lower perceived social status), gender (female vs. male), age (younger vs. older), fitness (physically favored vs. less fit individuals), and utilitarianism (smaller vs. larger group). Additionally, three secondary dimensions were incorporated: interventionism (swerving vs. continuing straight), relationship to the autonomous vehicle (passengers vs. pedestrians), and legal considerations (e.g., adherence to traffic signals). The scenarios were carefully constructed to balance complexity and clarity, ensuring that they effectively probed the ethical reasoning capabilities of the LLMs under study.
@@
To investigate how moral judgment patterns vary across different model architectures and evolve through updates, we expanded upon our previous study [34] by incorporating a comprehensive set of LLMs, including the latest versions available. Our analysis encompasses both base language models and models specifically optimized for dialogue interactions. While these models may differ in their training approaches, they share fundamental capabilities in processing and responding to complex queries, including moral scenarios. These LLMs were utilized to generate responses to moral machine scenarios.
@@
In the open-source domain, we included Meta’s Llama family [37], comprising Llama 3 and Llama 3.1 in 8B and 70B versions, Llama 3.2 in 1B and 3B versions, and Llama 3.3 in 70B version. A comprehensive evaluation of Google’s Gemma model family [38, 43] was conducted, encompassing multiple versions and scales from 2B-it to 27B-it, including the derivative DataGemma RIG 27B-it model [44]. Additional open-source models evaluated in this study included Large Model Systems Organization’s Vicuna v1.5 (7B and 13B) [45], Mistral (7B Instruct v0.2) [46], Mistral-NeMo [47], CohereForAI’s Command R+ [48], and Microsoft’s Phi-3.5 [49] (both MoE-instruct and mini-instruct versions).
@@
All models were used with their default parameter settings. The input prompts were consistent with those used in [34], requiring each model to select either “Case 1” or “Case 2” for each scenario. The complete code for prompt generation, analysis, and resulting data is available in our GitHub repository.
@@
To systematically evaluate how different LLMs align with human moral preferences and how this alignment changes across model versions and architectures, we analyzed the LLM responses using the conjoint analysis framework as described in [50] and implemented in the original study [11] on the Moral Machine experiment and our previous study [34]. This approach allows for a robust, non-parametric identification of causal effects without specific modeling assumptions.
@@ Results
Moral judgments of large language models To comprehensively evaluate how different LLMs approach moral decision-making in autonomous driving scenarios, we conducted a systematic analysis of their responses across multiple ethical dimensions. Our analysis framework focused on quantifying and comparing the moral preferences of various LLM families with human judgments.
Fig 1 presents radar charts characterizing the moral judgment tendencies of various LLMs across nine preferences quantified by Average Marginal Component Effect (AMCE) values (see S1 Table. for exact numerical data). These charts enable direct comparison of preferences across multiple versions and model sizes of major LLM families, including GPT-3.5, GPT-4, GPT-4o/o1, Claude, Gemini, Llama, Gemma, and other LLMs, against established human preferences (see S1 Fig. for detailed differences between each model and human values).
Our initial analysis revealed three key patterns in LLM moral judgments. First, many LLMs demonstrated strong alignment with human preferences in fundamental moral decisions, showing large positive values ($>0.5$) for No. Characters and Species dimensions. This indicates consistent tendencies in prioritizing the preservation of more lives and favoring humans over animals, suggesting that LLMs can capture basic human moral intuitions. Second, we observed significant variations in judgment patterns across different model families, with some models exhibiting inverse priorities compared to human preferences. Third, certain models showed disproportionately strong preferences in the same direction as humans (differences in AMCE values approximately $>0.3$ compared to human values), indicating potential overemphasis of specific moral principles.
These patterns provide important insights into both the capabilities and limitations of LLMs in moral reasoning tasks. Below, we elaborate on these findings for each model family, examining how different architectures influence moral judgments, and analyzing the effects of model updates on alignment with human preferences.
@@
Analysis of the GPT-3.5 family (Fig 1A) revealed significant temporal evolution in moral judgments across model updates. While models from June 2023 to January 2024 exhibited similar patterns, the March 2023 version showed distinctive tendencies, particularly in several key ethical dimensions.
In the Law dimension, the March version showed strong positive values favoring law-abiding individuals (AMCE = 0.7), but subsequent versions shifted toward more neutral judgments (AMCE $\approx$ 0). Similar temporal changes were observed in Gender-related judgments, where initial strong female preference (AMCE = 0.5 in March 2023) gradually reduced toward human judgment levels (AMCE = 0.2 in January 2024, human AMCE = 0.1). Fitness-related judgments also evolved from initially contradicting human preferences (AMCE $=-0.2$ in March 2023) to showing better alignment in later versions (AMCE = 0.1 in January 2024).
However, we observed persistent deviations from human preferences in certain dimensions. The latest versions maintained near-neutral age-related judgments (AMCE $\approx 0$) despite human preference for younger individuals (AMCE = 0.5), and showed consistently stronger preferences for pedestrians over passengers (AMCE = 0.6 vs. human 0.1). Species-related judgments demonstrated an interesting pattern where initial alignment with human values in earlier versions (AMCE = 0.7) weakened in later updates (AMCE = 0.2) while maintaining the same directional preference.
@@
Analysis of the GPT-4 family (Fig 1B) revealed remarkable consistency in moral judgments across different versions from March 2023 to January 2024. In contrast to GPT-3.5, model updates showed minimal impact on judgment patterns, with all versions maintaining similar preference strengths across ethical dimensions.
The GPT-4 family demonstrated strong alignment with human judgments in direction but often exceeded human preference magnitudes. Most notably, Species and No. Characters preferences showed near-deterministic values (AMCE $>0.9$) compared to human values (AMCE = 0.6), indicating substantially stronger preferences for saving humans over animals and larger groups over smaller ones. Similar amplification appeared in Gender (AMCE = 0.4 vs. human 0.1), Law (AMCE = 0.6 vs. human 0.3), and Relation to AV (AMCE = 0.5 vs. human 0.1) categories.
The only notable divergence appeared in the Fitness category, where GPT-4 models showed a slight preference for overweight individuals (AMCE $=-0.1$) contrary to human preference for physically fit individuals (AMCE = 0.2). This deviation, however, was minimal compared to the strong alignment observed in other dimensions.
@@
Analysis of the GPT-4o family (Fig 11) revealed a pattern of consistent moral judgments across models, with two notable exceptions: GPT-4o-mini and o1-mini showed distinctive characteristics. The family exhibited closest alignment with human preferences in Gender (AMCE = 0.1) and No. Characters (AMCE = 0.7) categories, matching human values within ±$\pm 0.1$ AMCE.
Most GPT-4o models showed stronger ethical preferences than both humans and the GPT-4 family in several dimensions. In Relation to AV, all models except GPT-4o-mini demonstrated extreme pedestrian preference (AMCE $>0.5$ vs. human 0.1). Similarly, they showed amplified preferences for law-abiding individuals (AMCE = 0.8 vs. human 0.4) and humans over pets (AMCE = 0.9 vs. human 0.6). However, these models showed weaker age-related preferences (AMCE = 0.2 vs. human 0.5) and maintained slight preferences for overweight individuals in Fitness judgments (AMCE $=-0.1$).
The o1 variants showed interesting divergences: while the standard o1 followed GPT-4o patterns, o1-mini demonstrated unique characteristics, including near-neutral Relation to AV preferences (AMCE = 0.1) but extremely strong preferences for larger groups (AMCE $\approx$ 1.0). This suggests that model size significantly influences moral judgment patterns within this family.
@@
Analysis of the Claude family (Fig 1D) revealed a clear bifurcation in moral judgment patterns, forming two distinct groups with contrasting ethical preferences. The first group (Claude 3 Sonnet, Opus, and 3.5 Haiku) and second group (Claude 3 Haiku and Claude 3.5 Sonnet) showed markedly different alignment with human preferences.
The first group demonstrated significant divergence from human judgments in several dimensions. They showed inverse preferences in Relation to AV (AMCE $=-0.4$), Fitness (AMCE $\approx-0.5$), and Social Status (AMCE $=-0.3$) categories, contrasting with human values (AMCE = 0.1, 0.2, and 0.3 respectively). While they aligned with humans in Gender and Species preferences, their magnitude was notably stronger in both dimensions (Gender: AMCE = 0.7 vs. human 0.3; Species: AMCE = 0.9 vs. human 0.6).
The second group showed different patterns: Claude 3 Haiku and Claude 3.5 Sonnet exhibited excessive preferences in Relation to AV (AMCE = 0.5) and Law (AMCE = 0.7) categories compared to humans (AMCE = 0.1 and 0.4, respectively). Claude 3.5 Sonnet particularly showed strong preferences in Species and No. Characters dimensions (AMCE $>0.8$).
Model updates showed inconsistent effects: while the Sonnet 3 to 3.5 update improved human alignment, the Haiku 3 to 3.5 update increased divergence. Subsequent updates within Claude 3.5 Sonnet showed minimal changes in judgment patterns.
@@
Analysis of the Gemini family (Fig 1E) revealed significant evolution in moral judgments across model updates, particularly marked by changes in the Relation to AV dimension. We observed a clear transition from Gemini 1.0 Pro to the 1.5 series, with distinct changes in ethical preferences.
The most notable shift occurred in the Relation to AV category, where Gemini 1.0 Pro initially showed preference for passengers (AMCE $=-0.4$), while the Gemini 1.5 Pro series shifted to strongly favor pedestrians (AMCE = 0.7 vs. human 0.1). The Gemini 1.5 Flash family showed a similar evolution: Flash Preview and 001 favored passengers (AMCE $=-0.7$), but Flash 002 shifted to strong pedestrian preference (AMCE = 0.7).
Gender and Species preferences also showed distinctive patterns. Only Gemini 1.5 Pro 002 aligned closely with human gender preferences (AMCE = 0.2 vs. human 0.1), while earlier versions showed stronger female preference (AMCE $>0.4$). All versions showed strong preferences than humans for prioritizing humans over pets (AMCE $>0.7$) and consistently showed inverse fitness-related preferences (AMCE $=-0.3$ vs. human 0.2).
Notably, compared to Google’s previous PaLM 2 model, which contradicted human preferences in group size decisions (AMCE = $-0.3$ vs. human 0.7), all Gemini versions maintained consistent alignment with human preferences in preserving larger groups (AMCE = 0.4–0.8).
@@
Analysis of the Llama family (Fig 1F) revealed that model size significantly influenced moral judgment patterns, with larger models showing more consistent alignment with human preferences. Most models demonstrated human-like preferences in the No. Characters category (AMCE = 0.2–0.9), with Llama 3.1 8B being the notable exception (AMCE $=-0.3$).
Ethical preferences varied substantially across different model sizes and versions. In gender-related decisions, smaller models (Llama 2 7B, 3.2 1B, 3.2 3B) showed male preference (AMCE $<-0.1$) contrary to human judgments (AMCE = 0.1), while larger models aligned with human preferences. Species-based decisions showed similar size-dependent patterns: 70B models (Llama 3, 3.1, and 3.3) demonstrated close alignment with human judgments (AMCE = 0.5–0.6), while smaller models (e.g., Llama 3.2 1B and 3B) showed inverse preferences (AMCE $<-0.1$).
The 70B models showed relatively small variations in judgments between versions and demonstrated closer alignment with human preferences. For example, these models maintained consistent preferences in law compliance (AMCE = 0.5) which happened to align with human values (0.4), while other models, excluding 3 8B model, exhibited neutral preferences. This observation suggests that model size may influence the consistency of judgment patterns.
@@
Analysis of the Gemma family (Fig 1G) revealed a mix of consistent patterns and model-dependent variations in moral judgments. Across all model variations, we observed consistent trends in law compliance (AMCE $\approx$ 0) and Relation to AV preferences (AMCE $<0$), with the latter diverging from human judgments (human AMCE = 0.1).
Judgment patterns varied significantly across different model sizes and versions. In the No. Characters category, smaller Gemma 2 models (2B, 9B) showed inverse preferences (AMCE $=-0.4$) compared to human values (AMCE = 0.7), while larger models aligned with human preferences. The Species category showed similar variation: the Gemma 2 series demonstrated human-aligned preferences (AMCE = 0.6), while other versions showed weak or neutral species preferences (AMCE $<0$). Social Status preferences varied notably, with only Gemma 7B and 1.1 7B matching human preferences (AMCE = 0.3), while other versions showed neutral or inverse preferences.
The variation in judgment patterns across different model sizes suggests that parameter count significantly influences moral reasoning capabilities, although this relationship appears more complex than in other model families.
@@
Analysis of other LLMs (Fig 1H) revealed distinct patterns of alignment and divergence from human moral preferences, with notable variations across model architectures. In the Gender category, most models except Vicuna showed substantially stronger female preference (AMCE $>0.4$) compared to human values (AMCE = 0.1).
Key differences emerged in utilitarian judgments: while Mistral and Command R+ aligned with humans in favoring larger groups, their preferences were markedly stronger (AMCE = 0.9 vs. human 0.6). Characteristic divergences appeared in the Relation to AV and Fitness categories, where most models except Vicuna 13B showed inverse preferences (Relation to AV: AMCE $<0$ vs. human 0.1; Fitness: AMCE $<0$ vs. human 0.2).
Vicuna models demonstrated unique patterns. While Vicuna 7B showed neutral judgments (AMCE $\approx$ 0) across multiple categories (Gender, Age, Species), Vicuna 13B exhibited distinctive preferences contrary to humans (Gender: AMCE = $-0.2$ vs. human 0.1; Species: AMCE = $-0.2$ vs. human 0.6). Similarly, Phi 3.5 models showed size-dependent variations, with the mini version demonstrating notably different preferences than the MoE version. These results suggest potential influences of model size on ethical judgments.
@@
Proprietary models generally showed closer alignment with human judgments than open-source alternatives. The GPT-4 family achieved the closest alignment (minimum distance = 0.6 with January 2024 version), followed by GPT-3.5 (distance = 0.8–0.9). However, the newer GPT-4o/o1 family showed slightly increased distances (0.9–1.0), with o1-mini being an exception (distance = 0.7).
Model updates showed varying effects across different families. In the Claude family, while Claude 3 Haiku and Claude 3.5 Sonnet demonstrated relatively close alignment with human judgments (distances = 0.9–1.0), Claude 3 Sonnet and Opus showed notably larger distances (1.2). The Gemini family similarly exhibited significant version-dependent variations, with Gemini 1.5 Preview (April 2024) showing the minimum distance of 0.9, while subsequent versions sometimes demonstrated larger distances of 1.1. Additionally, 1.5 Flash models showed greater distances compared to 1.5 Pro models.
Open-source model performance strongly correlated with model size. The Llama family showed clear size-dependent improvement, with 70B models achieving closer alignment (distance = 0.7–0.8) compared to smaller 1–8B models (distances = 1.2–1.6). Conversely, recent Gemma 2 models showed larger distances (2B/9B: distance = 2.0) despite being newer, though performance improved with size. Among other LLMs, Mistral-Nemo and Command R+ showed relatively good alignment (distances = 0.8–0.9), while Vicuna and Phi 3.5 demonstrated larger divergence (distances = 1.0–1.6).